# Constrained Language Generation with Discrete Diffusion Models

### v1

@@ Abstract
Constraints are critical in text generation as LLM outputs are often unreliable when it comes to ensuring generated outputs adhere to user defined instruction or general safety guidelines. To address this gap, we present Constrained Discrete Diffusion (CDD), a novel method for enforcing constraints on natural language by integrating discrete diffusion models with differentiable optimization. Unlike conventional text generators, which often rely on post-hoc filtering or model retraining for controllable generation, we propose imposing constraints directly into the discrete diffusion sampling process. We illustrate how this technique can be applied to satisfy a variety of natural language constraints, including (i) toxicity mitigation by preventing harmful content from emerging, (ii) character and sequence level lexical constraints, and (iii) novel molecule sequence generation with specific property adherence. Experimental results show that our constraint-aware procedure achieves high fidelity in meeting these requirements while preserving fluency and semantic coherence, outperforming auto-regressive and existing discrete diffusion approaches.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable advancements in reasoning, enabling progress across diverse domains. Beyond traditional NLP applications, LLMs are increasingly being integrated into decision-making pipelines in embodied agents, autonomous scientific systems, and industrial automation [1, 2, 3, 4]. Despite this promise, the stochastic nature of LLM outputs makes these models untrustworthy when required to adhere to even simple logical constraints, domain-specific knowledge, and safety-critical requirements [5].
This limitation manifests in well-documented issues such as hallucinations, misinformation, and toxic outputs, presenting severe risks in real-world deployments [6, 7]. Furthermore, generative language models remain susceptible to producing harmful or biased text, or even disregarding user instructions, particularly when confronted with adversarial or ambiguous prompts [8]. The inability of generative models to adhere to strict operational and safety constraints severely undermines their reliability in scientific and industrial applications. For instance, in autonomous laboratories, LLM-driven systems may misrepresent experimental constraints or respond to adversarial triggers, leading to erroneous instructions for synthesizing hazardous or unstable compounds [9, 10]. This lack of assurance raises concerns and diminishes trust in LLM-driven systems, limiting their adoption in safety-critical environments.
Recent efforts have attempted to improve constraint satisfaction in LLM outputs through soft alignment methods via reinforcement learning from human feedback (RLHF) [11], knowledge injection [12, 13], or heuristic filtering [14]. However, while these approaches reduce harmful outputs, they lack mechanisms to consistently provide conformance with domain-specific or complex logical requirements. Furthermore, techniques including post-processing and rejection sampling, either completely discard the generated content or significantly alter it, thus compromising the creative quality of the original generation.
@@
To address these challenges, we propose integrating token generation with a novel differentiable optimization process that directly guides outputs to meet the imposed constraints. A key enabling component is the recent development of discrete diffusion models [15, 16, 17, 18], which offer an alternative to conventional auto-regressive architectures and have demonstrated strong performance across language generation tasks [4, 19]. Analogous to continuous diffusion models, discrete diffusion progressively refines corrupted sequences into coherent outputs. Notably, these models have achieved perplexity scores comparable to similarly-sized auto-regressive architectures on recent benchmarks [19, 16]. In this work we show how combining discrete diffusion with structured constraints during the sampling process enables the imposition of domain or safety requirements at every denoising step, without sacrificing model creativity or expressiveness.
Our approach is tested with three representative use cases: (1) toxicity mitigation with tunable thresholds for safe LLM deployment, (2) instruction-following tasks requiring precise symbolic constraints, and (3) property adherence for molecular sequence generation. Experimental results demonstrate our method’s ability to obtain 0% threshold violations for toxicity, 100% satisfaction for instruction following constraints, and up to 3x valid and novel molecule generations.
Contributions. Our contributions are as follows:
@@
    We introduce CDD, a novel framework for constrained language generation that integrates discrete diffusion models with differentiable optimization, enforcing constraints directly within the diffusion sampling process.
@@
    To enable the constrained optimization process during the token generation, we formulate a projection based operator, which we implement using a Lagrangian Dual based framework.
@@
    We analyze CDD performance on a set of complex natural language and structured text generation tasks, reporting state-of-the-art results for both constraint satisfaction and qualitative metrics.
@@ 2 Related Work
Recent efforts in controllable text generation have largely focused on constrained decoding, where the output is dynamically restricted to adhere to predetermined syntactic, semantic, or lexical rules. Approaches in this area modify the decoding process to prune the vocabulary, ensuring that only tokens compliant with a formal grammar are considered [20], or employ external models to filter outputs when direct access to the model’s logits is limited [21]. Other methods have also adapted beam search to encourage the presence or absence of specific words [22, 23]. Although these techniques effectively guide the generation process, they depend on augmented sampling heuristics which encourage, but frequently fail to provide, satisfaction of even simple constraint sets. Therefore, a recent line of research has attempted at mapping natural language into formal logic representations to verify factual and logical consistency [24, 25, 26], and by enforcing feasibility constraints with external modules [27, 28]. However, these approaches, while effective in their respective domains, often require additional, domain-specific components to ensure that the generated outputs are not only syntactically valid but also adhere to practical or logical requirements. Most similar to our work, Kumar et al. [29] and Dathathri et al. [30] introduce gradient-based corrections into non-autoregressive and autoregressive language models, respectively, to steer text generation towards particular attributes; however, these methods suffer from notable short-comings, as outputs from Kumar et al. [29] exhibit poor performance (e.g., higher perplexity), and outputs from Dathathri et al. [30] often still violate imposed specifications. Our work draws on these insights, seeking to embed both syntactic and semantic constraints directly within the generative process for enhanced control and reliability.
@@ 3 Preliminaries
Generative diffusion models [31, 32] have revolutionized high-fidelity creation of complex data and serve as key components in state-of-the-art image and video synthesis pipelines [33, 34, 35]. Formally, they operate in two phases: a forward phase, which gradually introduces noise to clean data samples, followed by a reverse denoising phase, in which a deep neural network is trained to iteratively remove noise. The forward process defines a Markov chain $\{\bm{x}_{t}\}_{t=0}^{T}$, with initial sample $\bm{x}_{0}\sim p(\bm{x}_{0})$ and each transition $q(\bm{x}_{t}|\bm{x}_{t-1})$ obtained by adding Gaussian noise that increases as $t\to 1$. The reverse process relies on a learned network that directly predicts the noise added at each step, enabling the gradual recovery of the original data distribution. Once trained, this network can then be used to iteratively “denoise” random noise samples back into data-like samples. This step is also known as sampling.
@@
While diffusion models were originally developed for continuous data, they have recently been extended to discrete domains such as text, enabling non-autoregressive generation [36, 18, 16, 37]. In contrast to auto-regressive models which predict tokens one by one, discrete diffusion methods generate entire sequences in parallel by first corrupting sequences through a forward noising process and then iteratively reconstructing them with a learned reverse process. This is a key enabler recognized by this work, which exploits this modus operandi to impose global constraints while simultaneously maintaining high fidelity.
Let $\bm{x}_{0}=(\bm{x}_{0}^{1},\ldots,\bm{x}_{0}^{L})$ denote an input sequence of size $L$, where each token $\bm{x}_{0}^{i}\in\mathbb{R}^{V}$ is represented as a one-hot vector over a vocabulary of size $V$. In discrete diffusion models, the forward process produces, at time $t\in[0,T]$, a sequence of values $\bm{x}_{t}\in\mathbb{R}^{L\times V}$ that parametrizes probability distributions for each token $\bm{x}_{t}^{i}$ in the sequence. We denote the corresponding sequence of predicted tokens by $\bm{x}^{\star}_{t}=\arg\max(\bm{x}_{t})$ where the $\arg\max$ operator is applied to every member $\bm{x}_{t}^{i}$ of the sequence $\bm{x}_{t}$. The diffusion process specifies a forward transition, defined as:
@@
where $\alpha_{t}$ is a schedule that decreases over time, $\nu$ is a fixed reference distribution that specifies the type of corruption applied to each token, and $\mathrm{Cat}(\cdot;p)$ is the categorical distribution over probability vector $p\in\Delta^{V}$. For example, the Masked Diffusion Language Model (MDLM) [16] uses a one-hot vector corresponding to the [MASK] token. Here, once a token is replaced by [MASK] (i.e., absorbed), it remains fixed during the forward process, and once unmasked during the denoising process it remains unchanged. In contrast, Uniform Diffusion Language Models (UDLM) [17], uses $\nu$ as the uniform distribution over the vocabulary. Conversely, this instantiation allows tokens to be re-perturbed in later time steps.
4 Constraint-aware Language Generation
This section introduces our key contribution: a paradigm to provide constraint adherence during the token generation process. Our novel method, Constrained Discrete Diffusion (CDD) alternates differentiable projections with iterative denoising steps during the sampling phase, ensuring that generated sequences satisfy predefined rules.
4.1 Projected Diffusion Processes
Discrete diffusion models generate samples through a stochastic denoising process, wherein a sequence of discrete states $\bm{x}_{t}$ evolves toward a target distribution $p(\bm{x}_{0})$. At each iteration, the model updates the token representation via a transition function $\mathcal{U}_{\theta}$, parametrized by a learned weight vector $\theta$, producing the next-step estimate:
$\bm{x}_{s}=\mathcal{U}_{\theta}(\bm{x}_{t})$
where $s$ is the subsequent timestep in the continuous time reverse process and $0\leq s<t\leq T$. Here, $\mathcal{U}_{\theta}$ is a learned function that progressively refines $\bm{x}_{t}$ towards the target distribution by reversing the noising process defined by Equation (1).
@@ 4.1 Projected Diffusion Processes
While iteratively applying $\mathcal{U}_{\theta}$ enables sampling from the posterior distribution, it does not natively incorporate externally defined constraints, or even implicit ones, given the stochastic nature of the process. By introducing a constraint set $\mathbf{C}$ and enforcing $\bm{x}_{t}\in\mathbf{C}$ at each step, CDD extends the reverse process to sample from these constrained subdistributions. This is operationalized by the introduction of a projection operator:
$\bm{x}_{s}=\mathcal{P}_{\mathbf{C}}\bigl{(}\mathcal{U}_{\theta}(\bm{x}_{t})\bigr{)}$
@@
computing the “nearest sample” that falls within the feasible subdistribution according to a distance metric defined over probability distributions. CDD formulates this projection as a constrained optimization problem that minimizes the Kullback–Leibler (KL) divergence between the projected and original probability distributions, ensuring minimal deviation, while enforcing constraint adherence. Given the model’s predicted probabilities ${\bm{x}_{t}}\in\mathbb{R}^{L\times V}$, the projection is defined as:
@@
$\mathcal{P}_{\mathbf{C}}(\bm{x}_{t})=\arg\min_{\bm{y}}D_{\mathrm{KL}}\left(\mathcal{U}_{\theta}(\bm{x}_{t})\|\bm{y}\right)$ (2a)
[…]
where $\mathbf{C}$ represents the constraint set governing valid outputs. The projection step is formulated as a subproblem of the sampling procedure, which can be solved using gradient-based methods, as we discuss in section 4.2.
@@
Importantly, note that while the process relies on metrics computed at the token probability distributions, valid outputs are assessed at the sequence level. Indeed, the constraints are formulated such that the $\arg\max$ of the projected sequence $\bm{y}$ (referred to as $\bm{y}^{\star}$), must satisfy sentence level criteria $\bm{y}^{\star}\in\mathbf{C}$ [^1] We assume a greedy decoding scheme, hence the use of $\arg\max$, as is standard to current diffusion-based language models. However, this use of a non-differentiable $\arg\max$ operation poses challenges to the resolution of the projection, which relies on gradient-based optimization frameworks.
Differentiating over $\mathbf{\arg\max}$ operators.
@@
To address the non-differentiability of the $\arg\max$ function, we apply a Gumbel-Softmax relaxation $\tilde{\phi}$, which approximates discrete token probabilities with continuous values [38] as
$\tilde{\phi}(\bm{x}_{t}^{i})(v)=\frac{\exp\!\Bigl{(}\frac{\log\bm{x}_{t}^{i}(v)+g_{v}}{\text{T}_{\text{sample}}}\Bigr{)}}{\sum_{v^{\prime}=1}^{V}\exp\!\Bigl{(}\frac{\log\bm{x}_{t}^{i}(v^{\prime})+g_{v^{\prime}}}{\text{T}_{\text{sample}}}\Bigr{)}},$
@@
where $\bm{x}_{t}^{i}(v)$ is the probability of token $v$ in the vocabulary, $g_{v}$ is drawn from the Gumbel$(0,1)^{1}$ distribution for token $v$, and $\text{T}_{\text{sample}}>0$ is the temperature parameter controlling the smoothness of the output. This allows gradients to be propagated to solve the projection step while closely approximating the discretized $\arg\max$ operation.
4.2 Enforcing Constraints
@@
Consider a generic constraint on a sequence $\bm{x}$ defined via a measurable property given by a function $g(\bm{x})$. For instance, $g(\cdot)$ might evaluate sentence toxicity (see Section 5.1), count the number of words occurring in a string (see Section 5.2), or even serve as a black-box function computed by an external routine, as in our molecular sequence generation application (see Section 5.3). To guide the projection operations, we require a measure of constraint violation that can parametrize the update rule of the gradient-based projection method. For ease of presentation, we assume $g$ is non-negative and we express the constraint in the form $g(\bm{x})<\tau$, where $\tau\geq 0$ represents an acceptable threshold that must not be exceeded.
To quantify by how much a given sequence violates the constraint, we define,
$\Delta g(\tilde{\phi}(\bm{x}_{t}))=\max\left(0,g(\tilde{\phi}(\bm{x}_{t}))-\tau\right),$
where $\tilde{\phi}(\bm{x}_{t})$ is a projection or transformation that ensures $g$ is well-defined on $\bm{x}$. If $g(\tilde{\phi}(\bm{x}_{t}))\leq\tau$, the differentiable relaxation of the constraint is satisfied and $\Delta g$ is zero; otherwise, $\Delta g$ measures the violation.
@@
Since $\Delta g$ is typically non-linear and difficult to impose directly, we adopt a method inspired by the augmented Lagrangian approach [39], where the constraints are incorporated into the objective of a minimizer via Lagrange multipliers $\lambda$ and a quadratic penalty term $\mu$. Let $\mathcal{U}_{\theta}(\bm{x}_{t})$ be the probability distribution after the denoising step at diffusion time $t$. We introduce a projected distribution $\bm{y}$, which is iteratively updated to reduce the constraint violations (measured by the score $\Delta g$) while remaining close (in KL-divergence) to $\mathcal{U}_{\theta}(\bm{x}_{t})$. Concretely the agumented Lagrangian dual function is defined as:
@@
$\mathcal{L}_{\mathrm{ALM}}(\bm{y},\mathcal{U}_{\theta}(\bm{x}_{t});\lambda,\mu)=\;$ $D_{\mathrm{KL}}\left(\mathcal{U}_{\theta}(\bm{x}_{t})\|\bm{y}\right)$
+\;+ $\lambda\Delta g(\tilde{\phi}(\bm{y}))$
+\;+ $\frac{\mu}{2}\Delta g(\tilde{\phi}(\bm{y}))^{2}$

### v2

# Constrained Discrete Diffusion

@@ Abstract
Discrete diffusion models are a class of generative models that construct sequences by progressively denoising samples from a categorical noise distribution. Beyond their rapidly growing ability to generate coherent natural language, these models present a new and important opportunity to enforce sequence-level constraints, a capability that current autoregressive models cannot natively provide. This paper capitalizes on this opportunity by introducing Constrained Discrete Diffusion (CDD), a novel integration of differentiable constraint optimization within the diffusion process to ensure adherence to constraints, logic rules, or safety requirements for generated sequences. Unlike conventional text generators that often rely on post-hoc filtering or model retraining for controllable generation, CDD directly imposes constraints into the discrete diffusion sampling process, resulting in a training-free and effective approach. Experiments in toxicity-controlled text generation, property-constrained molecule design, and instruction-constrained text completion demonstrate that CDD achieves zero constraint violations in a diverse array of tasks while preserving fluency, novelty, and coherence while outperforming autoregressive and existing discrete diffusion approaches.
@@ 1 Introduction
Language generation, alongside many scientific problems, admits a natural representation as the generation of a discrete sequence from a finite alphabet. Examples range from natural language to molecular SMILES strings and linearized chemical procedures [1, 2, 3, 4]. While large language models (LLMs) and related sequence transformers have recently accelerated discovery by proposing candidate sequences with desirable attributes, their autoregressive sampling mechanism produces tokens sequentially, hindering the ability to provide a native mechanism to ensure constrained feasibility. When these constraints are not satisfied, the generated outputs may be unreliable and ineffective in real-world applications. For example, a single out-of-vocabulary atom in a SMILES string can render a synthesized compound meaningless or even dangerous; similarly an overlooked volume unit in an autonomous laboratory protocol can trigger unsafe reactions.
To limit such risks, LLMs are often deployed with a variety of guardrails. These include soft alignment through reinforcement learning from human feedback (RLHF), rejection sampling, or heuristic post-processing [5, 3]. However, these methods do not offer provable compliance with logical formulas, stoichiometric rules, or regulatory thresholds. This is exacerbated by the autoregressive nature of common language models, which generate sequences one token at a time, making it difficult to enforce constraints at the sequence level.
@@
In contrast, discrete diffusion models offer a compelling alternative generative mechanism [6, 7, 8, 9]. They refine a fully corrupted sequence by iteratively denoising the entire sample, and have shown strong performance across a variety of tasks, including text generation, molecular design, and program synthesis [4, 10, 7]. Since each step exposes a global view of the partial sequence, it creates a natural opportunity to impose sequence-level structure. This work introduces Constrained Discrete Diffusion (CDD), a framework that capitalizes on this property by coupling discrete diffusion with a differentiable projection operator. Given a corrupted sequence at a particular diffusion step, the projection searches for a new candidate that stays close to the model’s current score distribution, preserves entropy-based diversity, and simultaneously satisfies all user-defined constraints before the next denoising update. Because the introduced projection acts only at sampling time, the method is training-free and needs neither model retraining nor post hoc filtering. Its stochastic formulation also preserves the generative diversity, which is crucial for exploratory scientific design, as will be shown in the empirical analysis.
[…]
Contributions. This paper provides the following contributions:
@@
    It introduces Constrained Discrete Diffusion (CDD), a novel framework that integrates discrete diffusion models with a differentiable projection operator, enforcing global sequence-level constraints directly within the diffusion sampling process.
@@
    It formulates this projection operator by solving a Lagrangian dual program at each denoising step, making the constrained optimization tractable and effective for guiding sequence generation.
@@
    Through five extensive experiments on three different domains, we demonstrate that CDD achieves state-of-the-art constraint adherence (zero violations across all settings) while preserving high sample quality in terms of fluency, novelty, and coherence.
@@ 2 Related Work
Recent efforts in controllable text generation have largely focused on constrained decoding, where the output is dynamically restricted to adhere to predetermined syntactic, semantic, or lexical rules. Approaches in this area modify the decoding process to prune the vocabulary, ensuring that only tokens compliant with a formal grammar are considered [11], or employ external modules to filter outputs when direct access to the model’s logits is limited [12]. Other methods have also adapted beam search to encourage the presence or absence of specific words [13, 14]. Although these techniques effectively guide the generation process, they depend on augmented sampling heuristics which encourage, but frequently fail to provide, satisfaction of even simple constraint sets [15]. To address this, other research has implemented mappings from natural language to formal logic allowing verification of factual and logical consistency [16, 17, 18] and enforcement of feasibility constraints with external modules [19, 20]. However, these approaches, while effective in their respective domains, often require additional, domain-specific components to ensure that the generated outputs are not only syntactically valid but also adhere to practical or logical requirements.
Other controllable generation approaches rely on energy-based sampling procedures to impose constraints on the outputs. These methods typically define an energy function combining language model likelihood with constraint penalties and leverage algorithms such as Metropolis-Hastings to sample low-energy (i.e., high probability) sequences [21, 22, 23]. While such methods are highly effective for encouraging simple properties (e.g., style transfer), they inherit many of the theoretical and practical shortcomings of rejection sampling. In particular, hard constraints are typically enforced only probabilistically, making energy-based approaches less suitable for settings requiring strict constraint satisfaction, as is desirable for the tasks considered in this work.
Finally, several gradient-based sampling frameworks have been proposed to impose constraints on token sequences [24]. Building on this approach, Amini et al. [25] improve generation quality via structured Hamiltonian Monte Carlo. However, these methods lack mechanisms for enforcing hard constraints, and are limited to soft attribute control. Other gradient-based control methods for autoregressive generation, such as PPLM [26], apply conditional steering during decoding to bias generations toward desired attributes; we compare against this methods directly, and show that while it can shift output distributions, it fails to reliably enforce structural or semantic constraints.
@@ 3 Preliminaries
Generative diffusion models [27, 28] have revolutionized high-fidelity creation of complex data and serve as key components in state-of-the-art image and video synthesis pipelines [29, 30, 31]. Formally, they operate in two phases: a forward phase, which gradually introduces noise to clean data samples, followed by a reverse denoising phase, in which a deep neural network is trained to iteratively remove noise. The forward process defines a Markov chain $\{\bm{x}_{t}\}_{t=0}^{T}$, with initial sample $\bm{x}_{0}\sim p(\bm{x}_{0})$ and each transition $q(\bm{x}_{t}|\bm{x}_{s})$ obtained by adding Gaussian noise that increases as $t\to T$. Here, $s$ is the subsequent timestep in the continuous time diffusion process and $0\leq s<t\leq T$. The reverse process relies on a learned network that directly predicts the noise added at each step, enabling the gradual recovery of the original data distribution. Once trained, this network can then be used to iteratively “denoise” random noise samples back into data-like samples. This step is also known as sampling.
@@
While diffusion models were originally developed for continuous data, they have recently been extended to discrete domains such as text, enabling non-autoregressive generation [32, 9, 7, 33]. In contrast to autoregressive models which predict tokens one by one, discrete diffusion methods generate entire sequences in parallel by first corrupting sequences through a forward noising process and then iteratively reconstructing them with a learned reverse process. This is a key enabler recognized by this work, which exploits this modus operandi to impose global constraints while simultaneously maintaining high fidelity.
Let $\bm{x}_{0}=(\bm{x}_{0}^{1},\ldots,\bm{x}_{0}^{L})$ denote an input sequence of size $L$, where each token $\bm{x}_{0}^{i}\in\mathcal{V}$ is represented as a one-hot vector over a vocabulary $\mathcal{V}$ of $N$ distinct tokens. In discrete diffusion models, the forward process produces, at time $t\in[0,T]$, a sequence of values $\bm{x}_{t}\in\mathcal{V}^{L}$ that parametrizes probability distributions for each token $\bm{x}_{t}^{i}$ in the sequence. We denote the corresponding sequence of predicted tokens by $\bm{x}^{\star}_{t}=\arg\max(\bm{x}_{t})$ where the $\arg\max$ operator is applied to every member $\bm{x}_{t}^{i}$ of the sequence $\bm{x}_{t}$. The diffusion process specifies a forward transition, defined as:
@@
where $\alpha_{t}$ is a schedule that decreases over time, $\mathrm{Cat}(\cdot;p)$ is the categorical distribution over probability vector $p\in\Delta^{N}$, and $\nu$ is a fixed reference distribution that specifies the type of corruption applied to each token. For example, the Masked Diffusion Language Model (MDLM) [7] uses a one-hot vector corresponding to the [MASK] token for $\nu$. In contrast, Uniform Diffusion Language Models (UDLM) [8], uses $\nu$ as the uniform distribution over the vocabulary. Conversely, this instantiation allows tokens to be re-perturbed in later time steps.
In the reverse process, $\bm{x}_{T}$ is initialized from $\nu$ (e.g., MDLM initializes the probability vectors as one-hots where the weight is fully concentrated on the [MASK] token). For each position $i$ in the sequence, while $\bm{x}_{t}^{i}=\nu$, the denoiser parameterizes the transitions to clean data, $\bm{x}_{0}$, and the probability of $\bm{x}_{t}^{i}=\nu$ is discounted as $t\rightarrow 0$. Once $\bm{x}_{t}\neq\nu$, the probability vector transitions into a one-hot vector, concentrating all weight exclusively on the predicted token.
$q(\bm{x}_{s}\mid\bm{x}_{t},\bm{x}_{0})=\begin{cases}\mathrm{Cat}\bigl{(}\bm{x}_{s};\bm{x}_{t}\bigr{)},&\text{if }\bm{x}_{t}\neq\bm{\nu},\\[4.0pt] \mathrm{Cat}\!\Bigl{(}\bm{x}_{s};\frac{(1-\alpha_{s})\,\bm{\nu}+(\alpha_{s}-\alpha_{t})\mathbf{x}_{\theta}(\bm{x}_{t},t)}{1-\alpha_{t}}\,\Bigr{)},&\text{if }\bm{x}_{t}=\bm{\nu}.\end{cases}$ (2)
Equation (2) formalizes this reverse process parameterization, where $\mathbf{x}_{\theta}(\bm{x}_{t},t)$ denotes the trained denoiser.
4 Constraint Discrete Diffusion (CDD)
Projected diffusion sampling.
We begin by introducing a perspective of the discrete diffusion reverse process that motivates our approach. Prior work has shown that continuous score-based diffusion sampling processes can be framed as an sequential optimization procedure [34]. Score-based parameterizations enable this framing as the model learns to predict the gradients of the data density function, $\mathbf{x_{\theta}}(\bm{x}_{t},t)\approx\nabla\log q_{t}(\bm{x}_{t}\mid\bm{x}_{0})$ through Score Matching, and these gradients can then be applied to optimize $\bm{x}_{t}$ to the data distribution $q_{t}(\bm{x}_{t})$. This is typically conducted through a Langevin dynamics algorithms [35], which is used either directly [36] or through predictor-corrector Euler-discretizations [37]. In both cases, a series of $M$ steps of Langevin dynamics are taken for a fixed distribution $q_{t}(\bm{x}_{t})$:
$\bm{x}_{t(\ell)}^{\prime}\xleftarrow{\text{update~{}}(\ell)}\bm{x}_{t(\ell)}+\gamma_{t}\nabla_{\bm{x}_{t(\ell)}}\log q_{t}(\bm{x}_{t(\ell)}|\bm{x}_{0})+\sqrt{2\gamma_{t}}\bm{\epsilon}\qquad(\text{for }\ell=1\ldots M),$ (3)
where $\gamma_{t}$ is the step size, $\epsilon$ is added noise, and $\log q_{t}(\bm{x}_{t})$ is the density function of the learned distribution at time step $t$. Note that while annealing is used to improve convergence, it is applied only after every $M$ iterations. At that point, the sample is updated ($\bm{x}_{t(M)}\to\bm{x}_{t-1(1)}$) and the model transitions to the next distribution, $q_{t-1}(\bm{x}_{t-1}\mid\bm{x}_{0})$, however, the step size and the distribution $q_{t}(\bm{x}_{t}\mid\bm{x}_{0})$ remain stationary throughout these $M$ iterations.
Discrete diffusion models can leverage a discrete generalization of the score function, referred to as the Concrete score, to approximate the gradient of the probability density function [7, 9, 38]. Concrete Score Matching provides an approach mirroring continuous Score Matching in that the estimated gradients of the probability density function are used to guide the sample to high density regions of the target distribution. While not always explicitly framed as Concrete Score Matching [9, 38], the denoiser often implicitly models the score function, as supported by theoretical results in [7] demonstrating its equivalence in simplified formulations ($\nabla\log q_{t}(\bm{x}_{t}\mid\bm{x}_{0})\approx\langle\mathbf{x}_{\theta}(\bm{x}_{t},t),\mathbf{y}\rangle\;\;\forall\mathbf{y}=1,\ldots,N$, in this case). This enables the use of Langevin-based samplers for discrete diffusion, as commonly employed, e.g., in [38].
Effectively, under some regularity conditions, Langevin dynamics will converge towards a stationary point. As shown by Xu et al., Langevin dynamics acts as an “almost-minimizer” on the optimized function, which in this case will be the negative probability density function, converging within a fixed bound of the optimum. Hence, for each series of steps with a stationary density function, within this fixed bound, the sampling procedure can be framed as the optimization problem,
$\underset{{\bm{x_{t}}}}{\text{minimize}}\sum_{t=T}^{1}-\log q_{t}(\bm{x}_{t}|\bm{x}_{0})\;\;\;\text{s.t.}\;\;\bm{x}_{t}\in\mathbf{C}.$ (4)
@@ 4 Constraint Discrete Diffusion (CDD)
Here, Equation (4) extends the representation from an unconstrained optimization problem to a constrained version by introducing a constraint set $\mathbf{C}$. However, while iteratively applying the denoiser enables sampling from the posterior distribution, it does not natively incorporate externally defined constraints, or even implicit ones, given the stochastic nature of the process. Previous work for continuous modalities has proposed enforcing $\bm{x}_{t}\in\mathbf{C}$ by applying a Euclidean projection after each denoising step [34], which is natural for continuous modalities which fall in a real space, but this is misaligned when applied to discrete diffusion which operates over a probability simplex $\bm{x}_{t}\in\Delta^{N}$.
@@
To address this, we introduce a projection operator that minimizes the Kullback-Leibler (KL) divergence between the projected and original probability distributions, as opposed to a Euclidean distance metric. Given the model’s predicted probabilities, the projection is defined as:
@@
$\bm{x}_{s}=\bm{x}_{t(\ell+1)}=\mathcal{P}_{\mathbf{C}}\bigl{(}\bm{x}_{t(\ell)}^{\prime}\bigr{)}\stackrel{{\scriptstyle\text{def}}}{{=}}\arg\min_{\bm{y}}D_{\mathrm{KL}}\left(\bm{x}_{t(\ell)}^{\prime}\|\bm{y}\right)\;\;\;\text{s.t.}\;\arg\max(\bm{y})\in\mathbf{C}.$ (5)
The integration of this projection operator ensures that $\bm{x}_{s}$ is the “nearest sample” that falls within the feasible subdistribution according to a distance metric defined over probability distributions. This ensures the denoising trajectory remains within the allowable set when $\mathbf{C}$ is convex, enabling effective navigation along its boundary or interior toward an optimal solution (illustrated in Figure 2). Next, we show how this projection operator can be formulated as a subproblem within the sampling procedure and efficiently solved using gradient-based methods. Moreover, while convergence guarantees are available for convex constraint sets, as discussed below, Section 5 demonstrates how this technique can effectively handles highly non-linear constraints, including toxicity mitigation, molecular generation, and instruction following, achieving zero violations across all cases.
@@
Importantly, while state transitions are imposed on the token probability distributions, constraint satisfaction is evaluated on the decoded sequence. Indeed, the constraints are formulated such that the $\arg\max$ of the projected sequence $\bm{y}$ (referred to as $\bm{y}^{\star}$), must satisfy sentence level criteria $\bm{y}^{\star}\in\mathbf{C}$ [^1] We assume a greedy decoding scheme as is standard to current diffusion-based language models. However, this use of a non-differentiable $\arg\max$ operation also poses challenges to the resolution of the projection, which relies on gradient-based optimization.
@@
Differentiable projection. To address the non-differentiability of the $\arg\max$ function, we apply a Gumbel-Softmax relaxation $\tilde{\phi}$, which approximates discrete token probabilities with continuous values [40] as
$\tilde{\phi}(\bm{x}_{t}^{i})(v)=\frac{\exp\!\Bigl{(}\frac{\log\bm{x}_{t}^{i}(v)+\xi_{v}}{\text{T}_{\text{sample}}}\Bigr{)}}{\sum_{v^{\prime}=1}^{N}\exp\!\Bigl{(}\frac{\log\bm{x}_{t}^{i}(v^{\prime})+\xi_{v^{\prime}}}{\text{T}_{\text{sample}}}\Bigr{)}},$ (6)
@@
where $\bm{x}_{t}^{i}(v)$ is the probability of token $v$ in the vocabulary, $\xi_{v}$ is drawn from the Gumbel$(0,1)^{1}$ distribution for token $v$, and $\text{T}_{\text{sample}}>0$ is the temperature parameter controlling the smoothness of the output. This enables gradient propagation during the projection step, while closely approximating the discrete $\arg\max$ operation.
Augmented Lagrangian projection.
@@
Consider a generic constraint on a sequence $\bm{x}$ defined via a measurable property given by a function $g_{i}(\bm{x})$. For instance, $g_{i}(\cdot):\mathbb{R}^{L\times N}\!\to\!\mathbb{R}^{+}$ might evaluate sentence toxicity (see Section 5.1), count the number of words occurring in a string (see Section 5.3), or even serve as a black-box function computed by an external routine, as in our molecular sequence generation application (see Section 5.2). To guide the projection operations, we require a measure of constraint violation that can later inform the parametrization of our projection update rule; for ease of presentation, we express the constraint in the form $g_{i}(\bm{x})<\tau_{i}$, where $\tau_{i}\geq 0$ represents an acceptable threshold that must not be exceeded. To quantify by how much a given sequence violates the constraint, we define
$\Delta g_{i}(\tilde{\phi}(\bm{x}_{t}))=\max\left(0,g_{i}(\tilde{\phi}(\bm{x}_{t}))-\tau_{i}\right),$
where $g=(g_{1},\ldots,g_{m})$ can be treated as series of functions corresponding to a series of thresholds $\tau=(\tau_{1},\ldots,\tau_{m})$, and $\Delta g=(\Delta g_{1},\ldots,\Delta g_{m})$ is defined in analogously quantifying $m$ constraints.
@@
In practice, $\Delta g$ is non-linear, and, thus, to implement Equation (5), we adopt an augmented Lagrangian approach [41]. In augmented Lagrangian methods, the problem constraints are incorporated into the objective of a minimizer via Lagrange multipliers $\lambda$ and a quadratic penalty term $\mu$. Let $\bm{x}_{t}^{\prime}$ be the probability distribution after the denoising step at diffusion time $t$. We introduce a projected distribution $\bm{y}$, which is iteratively updated to reduce the constraint violations (measured by the score $\Delta g$) while remaining close (in KL-divergence) to $\bm{x}_{t}^{\prime}$. Concretely the augmented Lagrangian dual function is defined as:
@@
$\mathcal{L}_{\mathrm{ALM}}(\bm{y},\bm{x}_{t}^{\prime};\lambda,\mu)=\;$ $D_{\mathrm{KL}}\left(\bm{x}_{t}^{\prime}\|\bm{y}\right)+\sum_{i=1}^{m}\;\lambda_{i}\Delta g_{i}(\tilde{\phi}(\bm{y}))+\;\frac{\mu_{i}}{2}\Delta g_{i}(\tilde{\phi}(\bm{y}))^{2}$